OpenStudy (anonymous):

6 (-2x+1)= -3(4x+2) last equation I don't understand.

OpenStudy (anonymous):

Simply multiply out and rearrange. What problems are you having with it?

OpenStudy (anonymous):

Yes I know but when I do that I get a final answer of 6= -6 which is false. Is there no solution to the problem?

OpenStudy (anonymous):

you're correct.... if you simplify to a statement that "does not make sense", then there is no solution to that equation: 6 = -6 does not make sense so no solution.

OpenStudy (anonymous):

in this type of equations you should Simplify like this : \[-12x+6=-12x-6\] \[6=-6\] that's mean that the equation is imposible ......

OpenStudy (anonymous):

6 (-2x+1)= -3(4x+2) should expand like this -12+6=-12-6 is impossible